A General Power bond carries a coupon rate of 8.2%, has 9 years until maturity, and sells at a yield to maturity of 7.2%. (AsURGENT!!
Answer #1

1.
=1000*8.2%=82

2.
=1000*8.2%/7.2%*(1-1/1.072^9)+1000/1.072^9=1064.601613

3.
=1000*8.2%/6.2%*(1-1/1.062^9)+1000/1.062^9=1134.857572

Price increases by 70.25595933

4.
Current yield is more than yield to maturity